Find the Marginal Utility of X; the Marginal Utility of Y; the ratio MUx / MUy for each of the following:

a. U(x,y) = 15x.2y.3

b. U(x,y)=.5xy

c. U(x,y)=6x.4y.5

d. U(x,y)=3x.5y.8
Answer #1

a vax,y) = 15X. 2y. 3 MUX = au da = 1s () 27.3. = goy =au ay = 150.2(1). 3 =900 MUX MUR Вь хоь = R (b) u(x, y) = o.sxy MU = a

mum = .sy o. Sc C) Uх,9)= Gх, чу. 5 МО» - 5 — G ) - чЯ-S → ।२०५ М10 ч = 67 . u(O). с = 120 x их Muy = lroy (ао и а) 0 ( 9 ) =И04 - 350 . s(I) . 8 = 120X MUX 12 o , муу тэом